### v4

# Ricci-Notation Tensor Framework for Model-based Approaches to Imaging

Abstract
Model-based approaches to imaging, like specialized image enhancements in astronomy, facilitate explanations of relationships between observed inputs and computed outputs. These models may be expressed with extended matrix-vector (EMV) algebra, especially when they involve only scalars, vectors, and matrices, and with n-mode or index notations, when they involve multidimensional arrays, also called numeric tensors or, simply, tensors. While this paper features an example, inspired by exoplanet imaging, that employs tensors to reveal (inverse) 2D fast Fourier transforms in an image enhancement model, the work is actually about the tensor algebra and software, or tensor frameworks, available for model-based imaging. The paper proposes a Ricci-notation tensor (RT) framework, comprising a dual-variant index notation, with Einstein summation convention, and codesigned object-oriented software, called the RTToolbox for MATLAB. Extensions to Ricci notation offer novel representations for entrywise, pagewise, and broadcasting operations popular in EMV frameworks for imaging. Complementing the EMV algebra computable with MATLAB, the RTToolbox demonstrates programmatic and computational efficiency via careful design of numeric tensor and dual-variant index classes. Compared to its closest competitor, also a numeric tensor framework that uses index notation, the RT framework enables superior ways to model imaging problems and, thereby, to develop solutions.
I Introduction
“Tensors” easily associate with learning-based approaches to imaging. After all, TensorFlow from Google Brain [Abadi2016] and Tensor Comprehensions from Facebook AI [Vasilache2018] are machine learning systems connected with a well known deep learning revolution in object detection and classification. Nonetheless, imaging researchers like Mai et al. [Mai2022] highlight the importance of model-based tensor approaches, emphasizing explanatory value. Mai et al. also argue that a model-based approach helps to reduce the training data requirements for a learning-based approach, yielding what is in effect a hybrid approach.
Imaging researchers are finding advantages of model-based tensor approaches in a variety of applications. These include: multi-exposure, multi-focus, and hyper/multi-spectral image fusion to increase dynamic range, depth-of-field, and spatial resolution [Mai2022, Wang2023, Prevost2022]; inpainting, or tensor completion, of missing or corrupt pixels in images and videos, even rows or columns of the same [Liu2021, Bengua2017]; deblurring, or image restoration, posed as a constrained inverse problem often with added noise [Newman2020, Lefkimmiatis2015, Rezghi2017]; compressive sensing, a mixture of computational imaging and image compression, for light field imaging (LFI), synthetic aperature radar, and hyperspectral imaging modalities [Marquez2020, Qiu2020, Li2020]; other constrained image enhancements, like ones that enforce 3D spherical invariances [Skibbe2017]; and image quality assessments, like ones tailored for LFI [Zhou2020].
Model-based imaging research involving tensors [Mai2022][Zhou2020] exhibits patterns. Papers typically have a section or subsection, on fundamentals, that addresses unary, binary, and $N$-ary operations expressible with tensors, like multi-way or multi-index contractions, Kronecker and Khatri-Rao products, and Tucker, canonical polyadic (CP), and tensor trains (TT) constructs. More fundamental than operations, authors discuss an $n$-mode notation, especially for decomposition constructs, an index notation, also called Einstein notation, or a hybrid notation thereof. Sometimes, extended matrix-vector (EMV) operators are employed to represent Hadamard, or entrywise, and Kronecker, or outer, products. Notation can resemble MATLAB or Python+NumPy syntax for multidimensional arrays. Instead of matrix-vector (MV) or EMV algebra, authors prefer tensor algebra to represent and manipulate high-order image relationships, including generalizations of the singular value decomposition (SVD) for rank reduction and feature extraction, akin to principal components analysis, and optimizations where images are the unknown variables in a functional that suffers upon vectorization of variables.
When literature on learning-based [Abadi2016][Vasilache2018] and model-based [Mai2022][Zhou2020] approaches to imaging use the word, “tensor” mainly or solely means a numeric data structure, usually an MDA, together with numeric operations, like those identified with overloadable operators in object-oriented programming languages. This paper uses the word likewise, avoiding requirements for deeper tensor character addressed, for example by Synge and Schild [Synge1978], in books on geometric tensor calculus.
Whereas tensor notations and operations are in some senses implementation agnostic, a brief survey of tensor software is in order. Because its components, like the Cyclops Tensor Framework (CTF) [Solomonik2014, Singh2022], may apply to tensor contractions, completions, and decompositions for imaging, the NorthWest Chemistry (NWChem) software ecosystem is noteworthy [Valiev2010, Kowalski2021]. With the Tensor Contraction Engine (TCE) and the Tensor Algebra for Many-Body Methods (TAMM) [Baumgartner2005, Mutlu2023], this ecosystem targets massively parallel computing and large-scale molecular simulations. Intended more for imaging and for desktop parallel computing, Tensor Learning in Python (TensorLy) offers tensor decompositions to improve efficiency, robustness, and explainability of deep learning networks [Kossaifi2019, Panagakis2021]. For desktop computing with MATLAB, the Partial Least Squares (PLS) and Multivariate Image Analysis (MIA) Toolboxes, cited in biomedical imaging research [Cumpson2015, Bedia2020], offer tensor decompositions and imaging customizations.
As Harrison and Joseph argue [Harrison2016], the codesign of tensor algebra and software, what they call a tensor framework, has benefits. Codesign facilitates programmatic and computational efficiencies for solving a variety of model-based problems. In support, they cite Bader and Kolda’s [Bader2006, Kolda2009] development of $n$-mode+ notation and the MATLAB Tensor Toolbox (MTT), also called the Tensor Toolbox for MATLAB. While Bader and Kolda contributed notation for decomposition constructs along with original software, others leveraged existing notations while contributing decomposition constructs and MATLAB toolboxes. These include structured data fusion (SDF) with Tensorlab, led by Sorber and De Lathauwer [Sorber2015, Vervliet2014], and the TT-Toolbox with the TT construct, by Osledets [Oseledets2011].
This paper proposes the Ricci-notation tensor (RT) framework for model-based approaches to imaging. A successor to Harrison and Joseph’s numeric tensor (NT) framework [Harrison2016], itself a complement to an EMV framework, the RT framework inherits their advantages. The proposed framework comprises RT algebra and RT software, defined as the RTToolbox [Joseph2023], developed with this paper, plus MATLAB. The RT software parses the RT algebra, dispatching calculations to the MATLAB kernel, with no dependencies on other toolboxes.
Both the NT and proposed RT algebras are index notations that exploit and extend the Einstein summation convention. However, using a simpler dual-variant index notation, the RT algebra proves equally expressive as the NT algebra, shown here to be a multi-variant index notation, and aligns well with the Ricci notation of tensor calculus. Moreover, the RT algebra supports additional outer operations. Inspired by the broadcasting of EMV algebras, they remain grounded in a dual-variant index notation with summation convention.
The NT software had options [Harrison2016]. With LibNT, C/C++ end users expressed the NT algebra with programmatic and, especially, computational efficiency. With NTToolbox, comprising M-file and MEX wrappers for LibNT, MATLAB end users expressed the NT algebra with some programmatic and some computational penalties. The RT software has only the RTToolbox option [Joseph2023], although this paper elaborates on its design in a C/C++-compatible way. Comprising just M-files, no compilation or linking is required. As per MATLAB Central recommendations [MathWorks2022], avoidance of MEX files and binaries means that sharing is easy via File Exchange.
For its computational efficiency, the RT software leverages compiled unary and binary pagewise functions introduced, in 2020 and 2022, to the MATLAB kernel. Of note, the kernel pagewise multiplication [MathWorks2020] resembles the LibNT lattice multiplication [Harrison2016] introduced, in 2016, with the NT software. This paper elaborates on the lattice concept to facilitate RT left- and right-division using kernel functions. It also introduces pagewise concatenation, an interpreted function the RT software exploits to support $N$-ary tensor concatenation. Programmatic efficiency of the RT software is illustrated via software expressions cross-referenced to algebraic ones.
An image enhancement example, developed here, illustrates the RT algebra. Inspired by Sirbu et al.’s model-based research [Sirbu2013, Sirbu2016, Sirbu2017] for coronagraphy [Schulz2023], the example adopts a simple coronagraph model with two planes related by 2D discrete Fourier transforms and a phase aberration. Starting with a Creative Commons public domain (CC0) image [Wikipedia2022], called Airy_disk_D65, the impact of the model on a synthetic ground-truth image having an occulted star with nearby exoplanets is computed. The exoplanets become visible after correction of the aberration, treated as unknown.
The rest of this paper is organized as follows. Section II introduces the RT algebra, contrasting it with the NT algebra. Section III illustrates the RT algebra’s usefulness for a model-based approach involving 2D DFTs. Section IV summarizes the RT software, contrasting it with the NT software. Unlike Section II, Sections III and IV include MATLAB code. Finally, Section V highlights contributions of this paper, this time with reference to selected details from Sections II to IV.
II Tensor Algebra
In the proposed RT algebra, tensors are scalars, vectors, and matrices of any degree, where degree is the number of true-variant and/or false-variant indices. First, benefits of exploiting and extending Ricci notation for this purpose are introduced in the context of scalars, even though these benefits also apply to nonscalars. Second, additional extensions of the RT algebra are summarized in the context of vectors and matrices.
II-A Dual-Variant Index Notation
The proposed RT algebra is introduced with reference to the NT algebra. Using an underline operator in what its authors called a single-type index notation, the NT algebra exploited and extended Einstein summation to support arbitrary mixtures of inner, entrywise, and outer products of $N$-degree scalars, not called $N$-dimensional arrays with reason. Use of the word degree (or order), instead of dimensional, for the number of indices is consistent with a geometric tensor calculus.
In the NT algebra, the underline operator prevents Einstein summation, yielding entrywise products. One can demonstrate unnecessary complexity of the notation with simple examples. Consider the ternary inner product, $x$, and the ternary entrywise product, $y$, of three degree-one scalars, $a$, $b$, and $c$, with indices suppressed in inline math for readability:
$x$ $=a_{i}b_{i}c_{i}\text{,}$ (1)
$y_{i}$ $=a_{\underline{i}}b_{\underline{i}}c_{\underline{i}}\text{.}$ (2)
For a variety of derivations, one needs to pair operands of $N$-ary products arbitrarily. This is done with commutation and association identities, as follows for these examples:
$x$ $=b_{i}(a_{\underline{i}}c_{\underline{i}})\text{,}$ (3)
$y_{i}$ $=b_{\underline{i}}(a_{\underline{\underline{i}}}c_{\underline{\underline{i}}})\text{.}$ (4)
The underline operator stacks arbitrarily, meaning any index has a countable infinite number of variants. This freedom has implications for the NT software. It also means the NT algebra is actually a multi-type or multi-variant index notation.
Setting aside questions of contructivism, consider a simpler formalism, namely the RT algebra, that enjoys the accepted advantages of the NT algebra for numeric tensor purposes. With the RT algebra, inspired by the Ricci notation of geometric tensor calculus, any index always has two variants, true and false, that are identified usually with typeset position, subscript and superscript, respectively. The covariant and contravariant indices of the Ricci notation motivated these names.
With the Ricci notation, an Einstein summation applies to each index that repeats twice where one is covariant and the other is contravariant. Repeating indices of the same variant is disallowed for geometric reasons. This convention specifies inner products or, for one operand, contraction (summation) along each repeated index. With the RT algebra, indices of either variant need not be unique. Repeated indices of the same variant specify entrywise products or, for one operand, an attraction (selection) along each repeated index. Unique indices across multiple operands specify outer products.
To support $N$-ary expressions having any combination of inner, entrywise, and outer products, as opposed to a hierarchy of binary products indicated with parentheses or a left-to-right precedence rule, what matters for the RT algebra is whether all repeats of an index do not or do involve the same variant. Returning to the ternary inner and entrywise examples, they are expressed and paired as follows, respectively:
$x$ $=a_{i}b_{i}c^{i}\text{,}$ (5)
$=b_{i}(a^{i}c^{i})\text{,}$ (6)
$y_{i}$ $=a_{i}b_{i}c_{i}\text{,}$ (7)
$=b_{i}(a_{i}c_{i})\text{.}$ (8)
As with the NT algebra, the RT algebra enjoys full associativity and commutativity for $N$-ary products with scalar operands of arbitrary degree. With either algebra, the variant of one or more indices may change when altering associations.
Whereas inner products eliminate indices, tending to decrease the degree, outer products aggregate indices, tending to increase the degree. As they preserve repeated indices of the same variant, entrywise products tend to maintain the degree. Outer products are easily expressed as follows:
$z_{ijk}$ $=a_{i}b_{j}c_{k}\text{.}$ (9)
The examples of ternary inner, entrywise, and outer products are convenient for illustrating the known Ricci operation of contraction and an extended Ricci operation called attraction, both of which are unary operations that reduce degree:
$x$ $=z_{ii\bar{i}}\text{,}$ (10)
$y_{i}$ $=z_{iii}\text{.}$ (11)
An alternative way, namely an overscript bar, to specify a false-variant index is shown. Unlike with the underline of the NT algebra, each index of the RT algebra still has two possible variants. Odd or even stacked overscript bars collapse into one or zero overscript bars. To specify a contraction, at least one repeated index must have a complementary variant.
With a superscript/subscript notation, often preferred in literature on geometric tensor calculus, ambiguity is possible in expressions where not just an index but a tensor repeats. Consider the following apparently nonsensical statement:
$y_{i}^{j}$ $\neq y_{i}^{j}\text{.}$ (12)
With the overscript bar notation, one could instead write a sensible inequality to represent the asymmetry of a tensor:
$y_{i\bar{j}}$ $\neq y_{\bar{j}i}\text{.}$ (13)
Contrast this with the dot-spacer approach of Ricci notation, which allows superscripts for contravariant indices while clarifying index positions to resolve ambiguities as required:
$y_{i}^{.j}$ $\neq y_{.i}^{j}\text{.}$ (14)
The RT algebra allows the superscript/subscript notation, with dot spacers as required, and the proposed overscript bar notation, where all indices are subscripts and dot spacers are never required. Another case of relevance is when permutations matter. They generalize transpositions, even to scalars. Consider a permuted copy, $z$, of a degree-two scalar, $y$:
$z_{i}^{.j}$ $=y_{.i}^{j}\text{.}$ (15)
Removing the dot spacers here results in a sensical statement that, however, does not specify a numerical permutation:
$z_{i}^{j}$ $=y_{i}^{j}\text{.}$ (16)
The number of possible permutations grows with the factorial of degree. Context determines if they matter to a model.
Mixtures of inner, entrywise, and outer products, and ways in which to rearrange them, are significant to CP derivations like those used to illustrate the NT algebra. Consider the following key step of a degree-three CP derivation, which uses the RT overscript bar notation to clarify index positions and inner parentheses to group for asymptotic efficiency:
$u_{i\bar{\ell}}$ $=((v_{j\ell}v_{\bar{j}\ell^{\prime}})(w_{k\ell}w_{\bar{k}\ell^{\prime}}))\backslash(a_{ijk}v_{\bar{j}\ell^{\prime}}w_{\bar{k}\ell^{\prime}})\text{.}$ (17)
As with EMV algebras, a left (or right) division implies the solution to a linear system. In this case, it is as follows:
$(v_{j\ell}v_{\bar{j}\ell^{\prime}}w_{k\ell}w_{\bar{k}\ell^{\prime}})u_{i\bar{\ell}}$ $=a_{ijk}v_{\bar{j}\ell^{\prime}}w_{\bar{k}\ell^{\prime}}\text{.}$ (18)
Iterating the key step and others like it, a degree-three scalar, $a$, decomposes into three degree-two factors, $u$, $v$, and $w$:
$a_{ijk}=\sigma_{\ell}u_{i\bar{\ell}}v_{j\bar{\ell}}w_{k\bar{\ell}}\text{,}$ (19)
where the additional degree-one factor, $\sigma$, represents singular values, and where each degree-two factor obeys identities.
Previously, using the multi-variant index notation of the NT algebra, a novel derivation of the CP key step was provided. Because the same would be possible with the dual-variant index notation of the RT algebra, consider instead a difference between the NT and RT algebras concerning division, in the context of a simplified version of the CP key step:
[…]
$A_{\ell\ell^{\prime}}u_{i\bar{\ell}}$ $=b_{i\ell^{\prime}}\text{.}$ (21)
Provided that each index, in the denominator operand of a division, changes to its complement, the RT algebra is equally consistent as the NT algebra for implied linear systems.
With MATLAB’s implied EMV algebra, left/right division is equivalent to inversion or pseudo-inversion of the denominator operand followed by left/right multiplication. With the RT algebra, as with the NT algebra, there are multiple ways in which each index of a high-degree denominator operand may participate in a linear system, whether fully ranked or not. Absent the linear system context, fully specified by a left/right division, an inversion or pseudo-inversion of a denominator operand is ambiguous. Therefore, the RT algebra favours binary division expressions over a complicated scheme presented with the NT algebra to uniquely specify unary inverses.
In general, (partial) differentiation is explainable as the limiting solution of a linear system involving finite differences. The linear system may have an arbitrary mix of inner, entrywise, and outer products. Thus, (partial) differentiation rules for the RT algebra inherit observations made in the context of linear systems. To ensure consistency in formulations involving differentials, each index in the denominator operand of a differential changes to its complement when evaluated.
II-B Nonscalars and Broadcasting
Like the NT algebra, the RT algebra supports nonscalars having indices, i.e., vectors and matrices of nonzero degree. In this paper, scalars are typeset with a plain italics font and nonscalars with a bold Roman font. When the degree is nonzero, indices are indicated in display math. Indices are suppressed in inline math where possible due to context.
In general, the product of two nonscalars is noncommutative, which limits associativity for $N$-ary products. Consider a degree-two product, $\mathbf{C}$, of degree-two matrices, $\mathbf{A}$ and $\mathbf{B}$, where columns of the first operand undergo an inner product with rows of the second operand, and where indices of the operands express entrywise products as per variants:
$\mathbf{C}_{i\bar{j}}$ $=\mathbf{A}_{i\bar{j}}\mathbf{B}_{i\bar{j}}\text{.}$ (22)
As rows and columns, i.e., MV indices, may convert to or from tensor indices, there is always an equivalent expression, with scalar operands of higher degree, that is fully commutative and associative for arbitrary product expressions, e.g.:
$c_{sti\bar{j}}$ $=a_{s\bar{r}i\bar{j}}b_{rti\bar{j}}\text{,}$ (23)
$=b_{rti\bar{j}}a_{s\bar{r}i\bar{j}}\text{.}$ (24)
Here, first and second indices of scalar operands correspond to rows and columns, respectively, of nonscalar operands. For an $N$-ary product of degree-two scalars, the operand sequence may be rearranged arbitrarily. This is not possible, in general, for an equivalent $N$-ary product of degree-zero matrices.
The binary product is reused to illustrate how a contraction or attraction is replaceable with an inner or entrywise operation across two operands with attention to index variants:
$t$ $=\operatorname{tr}\mathbf{C}_{i\bar{i}}\text{,}$ (25)
$=\operatorname{tr}\mathbf{A}_{ii}\mathbf{B}_{\bar{i}\bar{i}}\text{,}$ (26)
$\mathbf{d}_{i}$ $=\operatorname{diag}\mathbf{C}_{ii}\text{,}$ (27)
$=\operatorname{diag}\mathbf{A}_{ii}\mathbf{B}_{ii}\text{.}$ (28)
Because of their resemblance to contraction and attraction, these examples include the trace operator, $\operatorname{tr}$, of MV algebra, and the main-diagonal operator, $\operatorname{diag}$, of EMV algebra, which concern only the rows and columns of an operand.
Inner products help justify why each enclosed index changes to its complement in a (conjugate) transposition. Consider these equivalent expressions for the Euclidean norm squared, $\|\mathbf{x}\|^{2}$, of an any-degree column vector, $\mathbf{x}$, having only false-variant indices, at first, and possibly-complex entries:
$\|\mathbf{x}^{\mathbf{k}}\|^{2}$ $=(\mathbf{x}^{\mathbf{k}})^{*}\mathbf{x}^{\mathbf{k}}\text{,}$ (29)
$=(\bar{\mathbf{x}}^{k_{1}k_{2}\dots k_{D}})^{\mathrm{T}}\mathbf{x}^{k_{1}k_{2}\dots k_{D}}\text{,}$ (30)
$=\bar{\mathbf{x}}_{k_{1}k_{2}\dots k_{D}}^{\mathrm{T}}\mathbf{x}^{k_{1}k_{2}\dots k_{D}}\text{,}$ (31)
$=\mathbf{x}_{\mathbf{k}}^{*}\mathbf{x}^{\mathbf{k}}\text{.}$ (32)
Here, indices of one variant in sequence are represented as an index vector. An overscript bar on a tensor, as opposed to an index, indicates complex conjugation of tensor entries.
Unlike the NT algebra, the RT algebra supports outer additions, subtractions, concatenations, and relations. The unifying idea is that any tensor of one degree may be expanded to an equivalent tensor of higher degree via an outer product with a unit tensor of the difference degree. Consider the following equations and relation, which consistently employ a degree-two matrix, $\mathbf{A}$, four degree-one column-vectors, $\mathbf{1}$, $\mathbf{b}$, $\mathbf{c}$, and $\mathbf{1}$, and an entrywise absolute-value operation, $|\cdot|$:
$\mathbf{A}_{i}^{j}$ $=\mathbf{1}_{i}(\mathbf{b}_{j})^{\mathrm{T}}+|\mathbf{c}_{i}|(\mathbf{1}_{j})^{\mathrm{T}}\text{,}$ (33)
$=\begin{bmatrix}\mathbf{1}_{i}&|\mathbf{c}_{i}|\end{bmatrix}\begin{bmatrix}(\mathbf{b}_{j})^{\mathrm{T}}\\ (\mathbf{1}_{j})^{\mathrm{T}}\end{bmatrix}\text{,}$ (34)
$\mathbf{A}_{i}^{j}$ $\geq\mathbf{1}_{i}(\mathbf{b}_{j})^{\mathrm{T}}\text{.}$ (35)
Whereas the NT algebra would require the with-unit outer products explicitly, as would Ricci notation, the RT algebra allows them implicitly, an extension to Ricci notation:
$\mathbf{A}_{i}^{j}$ $=(\mathbf{b}_{j})^{\mathrm{T}}+|\mathbf{c}_{i}|\text{,}$ (36)
$=\begin{bmatrix}\mathbf{1}&|\mathbf{c}_{i}|\end{bmatrix}\begin{bmatrix}(\mathbf{b}_{j})^{\mathrm{T}}\\ \mathbf{1}^{\mathrm{T}}\end{bmatrix}\text{,}$ (37)
$\mathbf{A}_{i}^{j}$ $\geq(\mathbf{b}_{j})^{\mathrm{T}}\text{.}$ (38)
The missing outer products are implied by what may be called natural algebra, setting aside index variant details.
For an easier-to-read example, this time with scalars, of the naturalness of the proposed RT algebra, consider a logarithmic model for the digital response or output, $y$, of a wide-dynamic-range image sensor to a uniform stimulus or input, $x$:
$y_{ij}$ $=a_{j}+b_{j}\ln(c_{j}+x_{i})+\epsilon_{ij}\text{.}$ (39)
Offset, gain, and bias parameters, $a$, $b$, and $c$, depend on the pixel index, $j$, but not on the illumination index, $i$. In addition to such structural noise, the model accounts for temporal and quantization noise, $\epsilon$. A parseable expression of RT algebra, the model implies three with-unit outer products, an entrywise function, and a mixed (entrywise-and-outer) product.
Outer RT operations resemble broadcasting, popularized by the implied EMV algebra of Python+NumPy, but are modelled with an index notation. Variants aside, one advantage of index-based outer operations is that they express broadcasting in a way that would look familiar to high-school students. This attests to the special value of the RT algebra for model-based approaches, which favour human intelligence, in relation to learning-based approaches, which favour machine intelligence. By design, the RT algebra includes and extends, with outer RT operations, the implied EMV algebra of MATLAB.
With MATLAB’s implied EMV algebra, an inner product of two matrices is representable as a product of vectors, with one conjugate transposed, after a vectorization is applied to each matrix. There is no MATLAB operator for an inner product of matrices, although there is one for an entrywise, or Hadamard, product of the same. In the RT algebra, these operators may be denoted with a bullet and a circle, respectively.
Although EMV algebra includes entrywise concatenation along non-row/column dimensions, computable in MATLAB via the kernel cat function, the RT algebra offers a way to model it with tensor indices, a model generalizable to outer concatenation. Consider an outer concatenation of degree-two matrices, $\mathbf{A}$ and $\mathbf{B}$, along a common tensor index:
$\mathbf{C}_{ijk}$ $=\operatorname{cat}_{j}\{\mathbf{A}_{ij},\mathbf{B}_{jk}\}\text{,}$ (40)
$=\delta_{j}^{\ell}(\alpha_{\ell}^{j}\mathbf{A}_{ij}+\beta_{\ell}^{j}\mathbf{B}_{jk})\text{.}$ (41)
The degree-three result, $\mathbf{C}$, is modelled with the help of degree-two scalars, $\alpha$ and $\beta$, each similar to a degree-two Dirac symbol, $\delta$, commonly used for index substitution:
$\alpha_{\ell}^{j}$ $=\begin{cases}1,&\ell=j\text{,}\\ 0,&\text{otherwise}\text{,}\end{cases}$ (42)
$\beta_{\ell}^{j}$ $=\begin{cases}1,&\ell=J_{1}+j\text{,}\\ 0,&\text{otherwise}\text{.}\end{cases}$ (43)
The constant, $J_{1}$, is the dimension size of the tensor index, $j$, in the first operand, $\mathbf{A}$. With additional constants, $J_{1}+J_{2}$, etc., the model is generalizable to $N$-ary concatenation.
III Imaging Example
After proposing a problem, inspired by an exoplanet imaging instrument called a coronagraph, this section applies the RT algebra and MATLAB to solve it. Results, presented first, summarize without equations or code a scalar optimization, called the phase aberration correction. Details, like a gradient matrix, are then modeled and simplified using (inverse) DFT operators and the RT algebra. Code, presented last, leverages (inverse) fast Fourier transforms for efficiency.
III-A Simulation Results
Figure 1 introduces a model-based example. It displays in false color, after each pixel value is squared, a ground-truth image synthesized from the grayscale version of a CC0-licensed source image, Airy_disk_D65. The latter simulates a focused spot of white light passing through a circular aperture and lens. Rings in the focal or image plane arise due to diffraction. The figure includes, with circles overlaid to define an annular mask, the source image in false color. Using the mask to model a coronagraph, a bright spot is blacked out, representing an occulted star. An outer symmetric region, where diffraction rings are faint, is likewise masked.
Figure 1: Source (left) and ground-truth (right) images (original scale). The $401\times 401$ pixel source is a CC0-licensed image, Airy_disk_D65 [Wikipedia2022], after conversion to grayscale and division by 255. The ground-truth image, shown after pixel values are squared, is the entrywise square root of the source image, after masking (occulting) symmetric inner and outer regions, defined by two circles (left overlays), and after adding spots to simulate twin exoplanets.
Given coronagraphy, the example uses an entrywise square root to make the ground-truth from the source, after conversion from class uint8 to double and division by 255. Both images have $401\times 401$ pixels. Filled circles introduced into the annular foreground simulate twin exoplanets in the ground-truth. Using bicubic and nearest-neighbour interpolation, respectively, ground-truth and mask images are downsampled or upsampled as needed, to vary dimension sizes equally for simulation purposes. The mask is of class logical.
The ground-truth image does not exhibit a phase aberration. To synthesize an aberrated image, a random phase offset is introduced entrywise in the Fourier domain, often called the pupil plane in the context of coronagraphy. Because pixel values in the image plane, the inverse Fourier domain, must be real, this phase aberration is constrained to obey the required symmetry. Notwithstanding a few entries, such as the DC value, each entry of the random offset, an image-like matrix, belongs to a uniform distribution of $\pi$ to $\pi$ radians.
Figure 2 addresses the main part, image enhancement, of the model-based example. It shows the input aberrated image and the output corrected image. Prior to display, each pixel value is squared. Starting with an initial guess of the phase aberration, a MATLAB optimization yields a corrected image, approaching the ground-truth, through iteration. The procedure updates the 2D aberration by minimizing a scalar sum squared error (SSE), which is overlaid. The SSE aggregates deviations in the image plane alone. A corrected image has nonnegative values in the annular foreground and zero values in the background.
Figure 2: Input (left) and output (right) images of an enhancement. A model-based approach addresses an unknown phase aberration in the pupil plane (Fourier domain) by minimizing an image-plane SSE. Nonzero background, and negative foreground, pixels define the SSE (right overlay). Diffraction rings and twin exoplanets become visible in the annular foreground when the 2D phase aberration is adequately corrected. Before display, pixel values are squared.
The initial guess of the phase aberration is a zero matrix. With it, the optimization converges toward a local minimum at which, although the phase aberration differs (not shown) from the initial random one, the problem is adequately solved, at least in terms of diffraction ring and exoplanet visibility.
In practice, a multistep computing problem is solveable only if it requires a reasonable amount of processing time, measurable in seconds ($\mathrm{s}$), and a reasonable amount of memory space, measurable in bytes ($\mathrm{B}$) of random access memory (RAM), for all steps. With desktop computing, reasonable space means the RAM of a personal computer (PC) suffices for the image resolution, $401\times 401$ pixels, at hand. Reasonable time means a fractional second per step suffices on the PC at hand, a Dell Latitude E7450 with $8\,\mathrm{GB}$ of RAM, for the same.
An accurate and efficient SSE software function is required for dimension sizes of interest, given a phase aberration, plus the same for first and second SSE derivatives with respect to the 2D aberration. Otherwise, the optimization may fail, due to numerical errors in approximations of the first derivative, the gradient, or may not succeed in reasonable time, due to higher-order curvature not captured without a second derivative, the Hessian, or second-derivative information, a Hessian multiply function (HMF). The RT algebra enables accurate and efficient SSE, gradient, and HMF functions in MATLAB.
Figures 3 and 4 report the compute time and memory space required, versus the number of pixels, by functions written to determine the SSE only, the SSE and the gradient, or the HMF only. Compute time, averaged over multiple invocations with a for loop, is measured using kernel tic and toc functions. Using kernel whos and sum functions, invoked with negligible overhead, the same MATLAB implementation totals and returns memory space reused by the written functions.
Figure 3: Compute time for imaging example (multiple scales). For an $M\times N$ problem, computing first (gradient) and second (HMF) derivatives of the scalar SSE, with respect to an $M\times N$ matrix (phase aberration in Fourier domain), is asymptotically equivalent to computing the SSE alone, thanks to time efficiencies of a model-based solution derived with the RT algebra.
Figure 4: Memory space for imaging example (multiple scales). For an $M\times N$ problem, computing first (gradient) and second (HMF) derivatives of the scalar SSE, with respect to an $M\times N$ matrix (phase aberration in Fourier domain), is asymptotically equivalent to computing the SSE alone, thanks to space efficiencies of a model-based solution derived with the RT algebra.
In MATLAB, the SSE, the gradient, and the HMF may be represented, without the RTToolbox, by a scalar, a 2D array, and a 3D array. Like the gradient, the HMF is not computed with every SSE computation. When computed, HMF time-and-space requirements per page are multiplied by the number of pages. This number, automatically chosen by an Optimization Toolbox routine, fminunc, appeared to be small enough, about two, for the SSE to be suitably minimized, as shown in Figure 2, in reasonable time and space overall.
III-B Scalar Error
The image enhancement input is an $M\times N$ aberrated image, $\mathbf{X}^{\text{a}}$. One defines a scalar SSE function in the image plane, an inverse Fourier domain, with a model. Corresponding to an SSE minimum, the image enhancement output is an $M\times N$ corrected image, $\mathbf{X}^{\text{t}}$, an estimate of the ground-truth. Figure 2 shows the images, $\mathbf{X}^{\text{a}}$ and $\mathbf{X}^{\text{t}}$, after entrywise squares.
To formulate the corrected image, initially unknown, one models an optical system having an entrywise $M\times N$ phase aberration, $\mathbf{\Phi}$, in the pupil plane, a Fourier domain:
$\mathbf{Y}^{\text{a}}$ $=\mathbf{U}\mathbf{X}^{\text{a}}\mathbf{V}^{\mathrm{T}}\text{,}$ (44)
$\mathbf{Y}^{\text{t}}$ $=\mathbf{Y}^{\text{a}}\circ\exp(\jmath\mathbf{\Phi})\text{,}$ (45)
$\mathbf{X}^{\text{t}}$ $=\frac{1}{MN}\bar{\mathbf{U}}^{\mathrm{T}}\mathbf{Y}^{\text{t}}\bar{\mathbf{V}}\text{.}$ (46)
Thus, given an aberration, $\mathbf{\Phi}$, the corrected image’s 2D DFT, $\mathbf{Y}^{\text{t}}$, is found from the 2D DFT, $\mathbf{Y}^{\text{a}}$, of the aberrated image, $\mathbf{X}^{\text{a}}$. An inverse 2D DFT yields a corrected image, $\mathbf{X}^{\text{t}}$.
The 1D DFT operators, $\mathbf{U}$ and $\mathbf{V}$, model an (inverse) 2D DFT via matrix multiplication. For an $M\times N$ matrix, $\mathbf{X}$, in the image plane and its $M\times N$ counterpart, $\mathbf{Y}$, in the pupil plane, the operators are matrices having dimension sizes $M\times M$ and $N\times N$, respectively. Each entry of the first operator, $\mathbf{U}$, is a complex number whose exponent varies with row and column positions of an integer outer product, $\mathbf{k}\mathbf{k}^{\mathrm{T}}$, modulo $M$:
$\mathbf{U}$ $=\exp(-\jmath 2\pi\mathbf{K}/M)\text{,}$ (47)
$\mathbf{K}$ $=\mathbf{k}\mathbf{k}^{\mathrm{T}}\mod M\text{,}$ (48)
$\mathbf{k}^{\mathrm{T}}$ $=\begin{bmatrix}0&\dots&M-1\end{bmatrix}\text{.}$ (49)
The second operator, $\mathbf{V}$, is defined similarly. The conjugate transpose of an operator gives its inverse, where a dimension size, $M$ or $N$, provides the scale needed for an identity:
$\frac{1}{M}\bar{\mathbf{U}}^{\mathrm{T}}\mathbf{U}$ $=\textbf{I}\text{.}$ (50)
By the problem definition, the aberrated image, $\mathbf{X}^{\text{a}}$, is real, and so is the corrected image, $\mathbf{X}^{\text{t}}$. Thus, the complex-valued 2D DFT, $\mathbf{Y}^{\text{a}}$, exhibits a phase symmetry. The 2D DFT of the corrected image, $\mathbf{Y}^{\text{t}}$, differs only in phase due to an offset. Therefore, for the corrected image to be real, there is an implied symmetry contraint on the phase aberration, $\mathbf{\Phi}$. Discarding the imaginary part, after an inverse 2D DFT, deals also with round-off error issues that arise in practice:
$\mathbf{X}^{\text{t}}$ $=\frac{1}{MN}\operatorname{Re}\{\bar{\mathbf{U}}^{\mathrm{T}}\mathbf{Y}^{\text{t}}\bar{\mathbf{V}}\}\text{.}$ (51)
To correct the aberrated image, one requires a priori information, in this case occultation. The corrected image, when occulted, defines background and foreground errors, $\mathbf{X}^{\text{b}}$ and $\mathbf{X}^{\text{f}}$, that are known to be zero in the ground-truth case:
$\mathbf{X}^{\text{b}}$ $=\mathbf{W}^{\text{b}}\circ\mathbf{X}^{\text{t}}\text{,}$ (52)
$\mathbf{X}^{\text{f}}$ $=\mathbf{W}^{\text{f}}\circ\mathbf{X}^{\text{t}}\text{,}$ (53)
$\mathbf{W}^{\text{f}}$ $=(\mathbf{1}-\mathbf{W}^{\text{b}})\circ u(-\mathbf{X}^{\text{t}})\text{.}$ (54)
The occultation mask, $\mathbf{W}^{\text{b}}$, is an input of the image enhancement. For the example, it is zero at each pixel inside an annulus and one outside it. Note the entrywise product model.
Background and foreground errors test for deviant pixel values, where entries of background and foreground masks, $\mathbf{W}^{\text{b}}$ and $\mathbf{W}^{\text{f}}$, are one, respectively. The masks are not complementary because a step function, $u$, filters out nonnegative values in the foreground. However, as the masks are orthogonal, they may be added to help formulate a single error image:
$\mathbf{X}^{\text{e}}$ $=(\mathbf{W}^{\text{b}}+\mathbf{W}^{\text{f}})\circ\mathbf{X}^{\text{t}}\text{.}$ (55)
If the error image, $\mathbf{X}^{\text{e}}$, deviates from a zero matrix then an aberration, $\mathbf{\Phi}$, requires correction. Thus, the SSE is defined as the inner product, $E$, of the error image with itself:
$E$ $=\bar{\mathbf{X}}^{\text{e}}\bullet\mathbf{X}^{\text{e}}\text{.}$ (56)
III-C Gradient Matrix
At a minimum of the SSE, its gradient, $\mathbf{\nabla}E$, with respect to the phase aberration, $\mathbf{\Phi}$, vanishes. To avoid differentiating a matrix, $\mathbf{X}^{\text{e}}$, with respect to a matrix, the model employs a degree-two scalar equivalent, $\phi$, of the phase aberration:
$\mathbf{\nabla}E$ $=(g^{ij}+\bar{g}^{ij})\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}\text{,}$ (57)
$g_{ij}$ $=\frac{\partial\bar{\mathbf{X}}^{\text{e}}}{\partial\phi^{ij}}\bullet\mathbf{X}^{\text{e}}\text{,}$ (58)
$\mathbf{\Phi}$ $=\phi^{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}\text{.}$ (59)
Degree-one basis operators, $\mathbf{e}$ and $\mathbf{e}^{\mathrm{T}}$, are column and row vectors, each with one nonzero, equal to one, whose vector index corresponds to the tensor index. When the error image, $\mathbf{X}^{\text{e}}$, is zeroed, the gradient of the SSE is also zeroed.
The gradient matrix, $\mathbf{\nabla}E$, appears to depend on an infinite-valued Dirac delta function, $\delta$, the derivative of a Heaviside step function, $u$, which is discontinuous at the origin. First, the partial derivative of the error image, $\mathbf{X}^{\text{e}}$, is expanded:
$\frac{\partial\mathbf{X}^{\text{e}}}{\partial\phi^{ij}}$ $=\frac{\partial\mathbf{W}}{\partial\phi^{ij}}\circ\mathbf{X}^{\text{t}}+\mathbf{W}\circ\frac{\partial\mathbf{X}^{\text{t}}}{\partial\phi^{ij}}\text{,}$ (60)
$\mathbf{W}$ $=\mathbf{W}^{\text{b}}+(\mathbf{1}-\mathbf{W}^{\text{b}})\circ u(-\mathbf{X}^{\text{t}})\text{,}$ (61)
where the sum, $\mathbf{W}$, of background and foreground masks, $\mathbf{W}^{\text{b}}$ and $\mathbf{W}^{\text{f}}$, is given a symbol. A problematic partial derivative term, which could be infinite valued, is then expanded:
$\frac{\partial\mathbf{W}}{\partial\phi^{ij}}\circ\mathbf{X}^{\text{t}}$ $=(\mathbf{1}-\mathbf{W}^{\text{b}})\circ\frac{\partial u(-\mathbf{X}^{\text{t}})}{\partial\phi^{ij}}\circ\mathbf{X}^{\text{t}}\text{,}$ (62)
$\frac{\partial u(-\mathbf{X}^{\text{t}})}{\partial\phi^{ij}}$ $=\delta(-\mathbf{X}^{\text{t}})\circ-\frac{\partial\mathbf{X}^{\text{t}}}{\partial\phi^{ij}}\text{.}$ (63)
Regrouping factors of the expanded derivative, via association and commutation properties of entrywise matrix products, an identity applies. It ensures a zero problematic derivative, via even and sifting properties of the Dirac delta function:
$\delta(\pm\mathbf{X}^{\text{t}})\circ\mathbf{X}^{\text{t}}$ $=\mathbf{0}\text{.}$ (64)
After eliminating the problematic derivative, a degree-two scalar, $g$, that defines the gradient, $\mathbf{\nabla}E$, is rewritten, thanks to an inner, entrywise, and outer product identity and the result, $\mathbf{W}$, of an implied entrywise mask product, $\mathbf{W}\circ\mathbf{W}$:
$g_{ij}$ $=\left(\mathbf{W}\circ\frac{\partial\bar{\mathbf{X}}^{\text{t}}}{\partial\phi^{ij}}\right)\bullet\mathbf{X}^{\text{e}}\text{,}$ (65)
$=\frac{\partial\bar{\mathbf{X}}^{\text{t}}}{\partial\phi^{ij}}\bullet(\mathbf{W}\circ\mathbf{X}^{\text{e}})\text{,}$ (66)
$=\frac{\partial\bar{\mathbf{X}}^{\text{t}}}{\partial\phi^{ij}}\bullet\mathbf{X}^{\text{e}}\text{.}$ (67)
Required for simplification, the inner, entrywise, and outer product identity is derived by considering an equivalent scalar expression that facilitates commutation and association:
$g_{ij}$ $=\left(w_{k\ell}\frac{\partial\bar{x}_{k\ell}^{\text{t}}}{\partial\phi^{ij}}\right)x_{\text{e}}^{k\ell}\text{,}$ (68)
$=\frac{\partial\bar{x}_{k\ell}^{\text{t}}}{\partial\phi^{ij}}(w^{k\ell}x_{\text{e}}^{k\ell})\text{.}$ (69)
Subsequently, using the inverse 2D DFT model, the partial derivative of the corrected image, $\mathbf{X}^{\text{t}}$, with respect to the degree-two scalar aberration, $\phi$, is expanded as follows:
$\frac{\partial\mathbf{X}^{\text{t}}}{\partial\phi^{ij}}$ $=\frac{1}{MN}\bar{\mathbf{U}}^{\mathrm{T}}\frac{\partial\mathbf{Y}^{\text{t}}}{\partial\phi^{ij}}\bar{\mathbf{V}}\text{,}$ (70)
$\frac{\partial\mathbf{Y}^{\text{t}}}{\partial\phi^{ij}}$ $=\mathbf{Y}^{\text{t}}\circ\jmath\frac{\partial\mathbf{\Phi}^{\text{t}}}{\partial\phi^{ij}}\text{,}$ (71)
$\frac{\partial\mathbf{\Phi}^{\text{t}}}{\partial\phi^{ij}}$ $=\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}\text{.}$ (72)
Similar to the nonconjugate case, the partial derivative of the corrected image’s conjugate, $\bar{\mathbf{X}}^{\text{t}}$, with respect to the degree-two scalar aberration, $\phi$, may be expressed as follows:
$\frac{\partial\bar{\mathbf{X}}^{\text{t}}}{\partial\phi^{ij}}$ $=\frac{-\jmath}{MN}\mathbf{U}^{\mathrm{T}}(\bar{\mathbf{Y}}^{\text{t}}\circ\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}})\mathbf{V}\text{.}$ (73)
To simplify this expression, as follows, the enclosed matrix, $\mathbf{Y}^{\text{t}}$, converts to an equivalent degree-two scalar, $y^{\text{t}}$:
$\mathbf{Y}^{\text{t}}\circ\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}$ $=y_{ij}^{\text{t}}\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}\text{.}$ (74)
As a scalar, of any degree, may be factored out of a multi-operand product, the partial derivative simplifies:
$\frac{\partial\bar{\mathbf{X}}^{\text{t}}}{\partial\phi^{ij}}$ $=\frac{-\jmath}{MN}\bar{y}_{ij}^{\text{t}}\mathbf{U}^{\mathrm{T}}\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}\mathbf{V}\text{.}$ (75)
Thus, the degree-two scalar, $g$, may be expressed in terms of another degree-two scalar, denoted $y^{\text{e}}$, as follows:
$g_{ij}$ $=\frac{-\jmath}{MN}\bar{y}_{ij}^{\text{t}}y_{ij}^{\text{e}}\text{,}$ (76)
$y_{ij}^{\text{e}}$ $=(\mathbf{U}^{\mathrm{T}}\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}\mathbf{V})\bullet\mathbf{X}^{\text{e}}\text{.}$ (77)
By replacing the basis operators, $\mathbf{e}$ and $\mathbf{e}^{\mathrm{T}}$, the 1D DFT operators, $\mathbf{U}$ and $\mathbf{V}$, and the error image, $\mathbf{X}^{\text{e}}$, with equivalent degree-two scalars, including Dirac symbols, $\delta$, the following reformulation arises, thanks to the RT algebra:
$y_{ij}^{\text{e}}$ $=(u_{\alpha k}\delta_{i}^{\alpha}\delta_{j}^{\beta}v_{\beta\ell})x_{\text{e}}^{k\ell}\text{,}$ (78)
$=u_{ik}x_{\text{e}}^{k\ell}v_{j\ell}\text{.}$ (79)
If each degree-two scalar, e.g., $x_{\text{e}}$, is now replaced with its equivalent matrix, i.e., $\mathbf{X}^{\text{e}}$, a hidden 2D DFT reveals:
$\mathbf{G}$ $=\frac{-\jmath}{MN}\bar{\mathbf{Y}}^{\text{t}}\circ\mathbf{Y}^{\text{e}}\text{,}$ (80)
$\mathbf{Y}^{\text{e}}$ $=\mathbf{U}\mathbf{X}^{\text{e}}\mathbf{V}^{\mathrm{T}}\text{.}$ (81)
Consequently, the matrix equivalent, $\mathbf{G}$, of the degree-two scalar, $g$, is directly proportional to an entrywise product, in the Fourier domain with one operand conjugated, of the corrected image, $\mathbf{X}^{\text{t}}$, and the error image, $\mathbf{X}^{\text{e}}$. As the gradient matrix, $\mathbf{\nabla}E$, equals twice the real part, $\operatorname{Re}\{\mathbf{G}\}$, of this equivalent, its final formulation proves to be relatively simple:
$\mathbf{\nabla}E$ $=\mathbf{G}+\bar{\mathbf{G}}\text{,}$ (82)
$=\frac{2}{MN}\operatorname{Im}\{\bar{\mathbf{Y}}^{\text{t}}\circ\mathbf{Y}^{\text{e}}\}\text{.}$ (83)
III-D Hessian Function
For the phase aberration correction problem, once the SSE and its gradient are formulated, one may attempt to construct a solution using an established routine, like fminunc from MATLAB’s Optimization Toolbox. This function offers the user a choice between quasi-Newton and trust-region algorithms. Because the former proves too slow even on small-format test cases, all simulations employ the latter. With the quasi-Newton algorithm, the gradient is actually optional. With the trust-region algorithm, the gradient is required.
Asymptotically, it takes equal time and space to construct the SSE and gradient as the SSE alone. Using explicit DFT operators, either takes $O(M^{2}N+MN^{2})$ time and $O(MN)$ space. Each DFT operator represents a Fourier transform, or its inverse, as matrix multiplication. One need not construct the operator, a useful formalism, as the FFT constructs a DFT faster. Using 2D FFTs, the SSE and gradient or the SSE only takes $O(MN\log MN)$ time and $O(MN)$ space.
Non-asymptotically, it takes longer to compute the SSE and gradient than it does to compute the SSE alone. Yet without the gradient, on small-format test cases, the trust-region algorithm was too slow. For medium and large-format test cases, even the benefit of an explicit first-order partial derivative, the gradient, was insufficient. For a practical solution, an implicit second-order partial derivative, an HMF, was needed. The RT algebra helps to formulate and simplify the desired HMF.
On one hand, the Hessian, $\mathbf{H}$, of the scalar SSE, $E$, is a degree-two matrix that requires at least $O(M^{2}N^{2})$ space and, therefore, time to construct, even with FFT acceleration:
$\mathbf{H}_{ij}$ $=\frac{\partial\mathbf{\nabla}E}{\partial\phi^{ij}}\text{.}$ (84)
On the other hand, the HMF, $\mathbf{F}$, is a degree-one matrix that requires at least $O(MNP)$ space and time to construct, where one dimension size, $P$, refers to a third index, $k$, of a phase step, $\Delta\phi$, a degree-three, not a degree-two, scalar here:
$\mathbf{F}_{k}$ $=\mathbf{H}_{ij}\Delta\phi_{k}^{ij}\text{.}$ (85)
According to details of the trust-region algorithm, the HMF is computed occasionally only after a gradient computation. Referring to 2D DFTs, $\mathbf{Y}^{\text{t}}$ and $\mathbf{Y}^{\text{e}}$, of corrected and error images, $\mathbf{X}^{\text{t}}$ and $\mathbf{X}^{\text{e}}$, the HMF is readily expanded:
$\mathbf{F}_{k}$ $=\frac{2}{MN}\operatorname{Im}\left\{\frac{\partial\bar{\mathbf{Y}}^{\text{t}}}{\partial\phi^{ij}}\circ\mathbf{Y}^{\text{e}}+\bar{\mathbf{Y}}^{\text{t}}\circ\frac{\partial\mathbf{Y}^{\text{e}}}{\partial\phi^{ij}}\right\}\Delta\phi_{k}^{ij}\text{,}$ (86)
$=\frac{2}{MN}\operatorname{Im}\{\Delta\bar{\mathbf{Y}}_{k}^{\text{t}}\circ\mathbf{Y}^{\text{e}}+\bar{\mathbf{Y}}^{\text{t}}\circ\Delta\mathbf{Y}_{k}^{\text{e}}\}\text{,}$ (87)
where HMF components, $\Delta\bar{\mathbf{Y}}^{\text{t}}$ and $\Delta\mathbf{Y}^{\text{e}}$, represent a sequence of 2D DFT increments, with one of them conjugated. The first HMF component, $\Delta\mathbf{Y}^{\text{t}}$, is also readily formulated:
$\Delta\mathbf{Y}_{k}^{\text{t}}$ $=(\mathbf{Y}^{\text{t}}\circ\jmath\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}})\Delta\phi_{k}^{ij}\text{,}$ (88)
$=\jmath\mathbf{Y}^{\text{t}}\circ\Delta\mathbf{\Phi}_{k}\text{,}$ (89)
$\Delta\mathbf{\Phi}_{k}$ $=\Delta\phi_{k}^{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}}\text{,}$ (90)
where the degree-three scalar, $\Delta\phi$, is replaced with an equivalent degree-one matrix, $\Delta\mathbf{\Phi}$. The incremental time and space cost of constructing this conversion, if it were actually needed beyond a formalism, can be made practically zero.
The second HMF component, $\Delta\mathbf{Y}^{\text{e}}$, is partially formulated by revealing a sequence of 2D DFTs as follows:
$\Delta\mathbf{Y}_{k}^{\text{e}}$ $=\left(\mathbf{U}\frac{\partial\mathbf{X}^{\text{e}}}{\partial\phi^{ij}}\mathbf{V}^{\mathrm{T}}\right)\Delta\phi_{k}^{ij}\text{,}$ (91)
$=\mathbf{U}\Delta\mathbf{X}_{k}^{\text{e}}\mathbf{V}^{\mathrm{T}}\text{,}$ (92)
$\Delta\mathbf{X}_{k}^{\text{e}}$ $=\left(\mathbf{W}\circ\frac{\partial\mathbf{X}^{\text{t}}}{\partial\phi^{ij}}\right)\Delta\phi_{k}^{ij}\text{,}$ (93)
$=\mathbf{W}\circ\Delta\mathbf{X}_{k}^{\text{t}}\text{.}$ (94)
To complete the formulation, a sequence of inverse 2D DFTs of the first HMF component, $\Delta\mathbf{Y}_{k}^{\text{t}}$, is revealed:
$\Delta\mathbf{X}_{k}^{\text{t}}$ $=\frac{\jmath}{MN}\bar{\mathbf{U}}^{\mathrm{T}}(\mathbf{Y}^{\text{t}}\circ\mathbf{e}_{i}\mathbf{e}_{j}^{\mathrm{T}})\bar{\mathbf{V}}\,\Delta\phi_{k}^{ij}\text{,}$ (95)
$=\frac{\jmath}{MN}\bar{\mathbf{U}}^{\mathrm{T}}(\mathbf{Y}^{\text{t}}\circ\Delta\mathbf{\Phi}_{k})\bar{\mathbf{V}}\text{,}$ (96)
$=\frac{1}{MN}\bar{\mathbf{U}}^{\mathrm{T}}\Delta\mathbf{Y}_{k}^{\text{t}}\bar{\mathbf{V}}\text{.}$ (97)
Thanks to the RT algebra, the HMF asymptotically requires, with 2D FFTs, $O(MNP\log MN)$ time and $O(MNP)$ space to compute. This asymptotic complexity is directly proportional by a factor, $P$, to that of the gradient. The incremental time-and-space complexity of computing the HMF, $\mathbf{F}$, with the gradient, $\mathbf{\nabla}E$, is $O(1)$ asymptotically if $P$, which depends on the optimization routine, fminunc, is bounded above.
III-E Implementation
Figures 5 and 6 summarize the phase aberration correction solution. These figures give MATLAB functions to compute the SSE, its gradient, and its HMF, with software statements cross-referenced to algebraic equations. The code also shows how to compute a corrected image, $\mathbf{X}^{\text{t}}$, from the aberrated image, $\mathbf{X}^{\text{a}}$, and an optimizable phase aberration, $\mathbf{\Phi}$.
function [E,GrE,Xt] = ssefun(Ph,Xa,Wb)
Ya = fft2(Xa); % (44)
Yt = Ya.*complex(cos(Ph),sin(Ph)); % (45)
Xt = real(ifft2(Yt)); % (51)
Wf = ˜Wb & Xt < 0; % (54)
Xe = (Wb | Wf).*Xt; % (55)
if nargout > 1
      MN = numel(Xt);
      Ye = fft2(Xe); % (81)
      GrE = 2/MN*imag(conj(Yt).*Ye); % (83)
end
Xe = Xe(:);
E = Xe’*Xe; % (56)
Figure 5: Model-based SSE and gradient. Developed for the imaging example, this MATLAB function’s output arguments are the SSE, $E$, the SSE gradient, $\mathbf{\nabla}E$, and a corrected image, $\mathbf{X}^{\text{t}}$. Its input arguments are a phase aberration, $\mathbf{\Phi}$, the aberrated image, $\mathbf{X}^{\text{a}}$, and the occultation (background) mask, $\mathbf{W}^{\text{b}}$.
function F = hessmfun(Xt,DPh,Wb)
Yt = fft2(Xt); % (45)
W = Wb | (˜Wb & Xt < 0); % (61)
Ye = fft2(W.*Xt); % (55) & (81)
[M,N] = size(Xt);
DPh = reshape(DPh,M,N,[]);
MN = numel(Xt);
DYt = 1j*(Yt.*DPh); % (89)
DXt = real(ifft2(DYt)); % (97)
DYe = fft2(W.*DXt); % (94) & (92)
DYY = imag(conj(DYt).*Ye); % (87)
YDY = imag(conj(Yt).*DYe); % (87)
F = 2/MN*(DYY+YDY); % (87)
F = reshape(F,MN,P);
Figure 6: Model-based multipage HMF. Invoked occasionally after another function, ssefun, this MATLAB function’s output argument is the HMF, $\mathbf{F}$. Its input arguments are a corrected image, $\mathbf{X}^{\text{t}}$, a multipage phase aberration step, $\Delta\mathbf{\Phi}$, and the occulatation mask, $\mathbf{W}^{\text{b}}$. Some lines use broadcasting.
The MATLAB functions, ssefun and hessmfun, do not require the RTToolbox. However, they exploit efficiencies derived using the RT algebra. Figures 5 and 6 clarify the (inverse) 2D DFTs revealed after algebraic manipulations that involved purely numeric aspects of tensor operands, some nonzero degree. The dual-variant index notation excelled on a variety of inner, entrywise, outer, and mixed products.
Where one of two operands is multipage, a 3D array, the HMF, hessmfun, leverages EMV broadcasting. An operator, .*, enables entrywise products with pagewise broadcasting. The HMF uses pagewise kernel routines, fft2 and ifft2, also. Thus, the example motivates nonscalar and broadcasting extensions of Ricci notation featured by the RT algebra.
IV Tensor Software
Table I introduces the RTToolbox. Together with the classes and functions of the MATLAB kernel, compiled or interpreted, the RTToolbox defines the RT software. First, this section offers a design overview of tensor objects. Next, it presents index objects, which play a critical role in parsing a dual-variant index notation. Finally, key problems and solutions, contrasted with closest equivalents of the NT software, are summarized for unary, binary, and $N$-ary operations.
TABLE I: Composition of the RTToolbox. The RT software is defined as the RTToolbox, comprising tensor and index classes, pagewise functions, and unit tests, and the MATLAB kernel. Together, they naturally express and compute the RT algebra. There are no dependencies on other toolboxes (or libraries).
M-file Summary Note
tensor.m Definition of tensor objects Table II
tensorTest.m Tests of tensor objects Table III
index.m Definition of index objects Table IV
indexTest.m Tests of index objects Table V
pagetrace.m Pagewise trace function Figure 7
pagediag.m Pagewise diag function -
pagehorzcat.m Pagewise horzcat function Figure 7
pagevertcat.m Pagewise vertcat function -
pagecat.m Pagewise cat function -
IV-A Tensor Objects
Table II summarizes the tensor class. Apart from the first group, constructor aside, and isequal, all methods return a tensor object. Table III presents excerpts, cross-referenced to selected equations, from unit tests of class tensor.
TABLE II: Tensor object class definition. Favouring simplicity, this release of the tensor class has a constructor and just these public methods, including overloaded operators. Ordinary methods are grouped to further simplify the discussion.
Public methods (tensor class)
Constructor, etc.
    tensor (0, 1, 2+ args in), index, entry, degree, ndims,
    numel, length, size, end ((,))
Unary operations
    subsref ((,)), sum, permute, uminus (-), uplus (+),
    conj, not (~), abs, log, round, transpose (.’),
    ctranspose (), trace, diag
Binary operations
    subsasgn ((,)=), plus (+), minus (-), eq (==), ne (~=),
    lt (<), gt (>), le (<=), ge (>=), and (&), or (|),
    times (.*), ldivide (.\), rdivide (./), power (.^),
    mtimes (*), mldivide (\), mrdivide (/)
NN-ary operations
    horzcat ([,]), vertcat ([;]), cat, isequal
TABLE III: Tensor object unit tests. Included with the RTToolbox, the tensor object unit tests were developed for this paper. To illustrate programmatic efficiency, software expressions are cross-referenced to equivalent algebraic expressions. Here, DN means any degree and D# means degree #.
Algebraic expression (ref.) Software expression
Inner product, D1 scalar (5) a(i)*b(i)*c(~i)
Entrywise product, D1 scalar (7) a(i)*b(i)*c(i)
Outer product, D1 scalar (9) a(i)*b(j)*c(k)
Entrywise relation, D2 scalar (13) y(i,~j) ~= y(~j,i))
Permute and copy, D2 scalar (15) z(i,~j) = y(~j,i)
[…]
Mixed product, D2 scalar (21) A(l,lp)*u(i,~l)
Outer addition, D1 scalar (39) log(c(j)+x(i))
Mixed product, D2 matrix (22) A(i,~j)*B(i,~j)
Trace contraction, D2 matrix (25) trace(C(i,~i))
Diagonal attraction, D2 matrix (27) diag(C(i,i))
Inner product, DN vector (29) x(~k)’*x(~k)
Col. concatenation, D1 vector (37) [ones(M,1) abs(c(i))]
Row concatenation, D1 vector (37) [b(j).’; ones(1,N)]
Outer relation, D1 vector (38) A(i,~j) >= b(j).’
Index concat., D2 matrix (40) cat(j,A(i,j),B(j,k))
A tensor object has two protected properties, indices and entries, which the index and entry methods offer a copy of, respectively. The indices property is a row vector, possibly empty, of class index. The entries property must be a non-tensor array, an MDA, usually of class double or logical, the only classes represented in unit tests.
For a tensor object, the degree method returns the number of indices, i.e., the numel of the indices property, and the ndims method returns the number of dimensions, i.e., the ndims of the entries property. Because first and second dimensions are MDA rows and columns, the degree is at least two less than the ndims of a tensor object.
The one-argument constructor returns a tensor object, whose entries and degree equal the non-tensor argument and its ndims minus two. A corresponding index object is constructed and assigned to the indices. A zero-argument, or default, constructor returns a tensor object of degree zero with empty entries and indices.
The last constructor form enables a tensor object of specifiable indices and degree, in addition to entries. A first non-tensor argument is assigned to the entries. Additional arguments must be index scalars or row vectors. The concatenated index arguments, extended if the result is too short compared to ndims minus two of the entries, is assigned to the indices. The degree exceeds ndims minus two when index arguments suffice to address trailing singleton (size one) dimensions of the entries.
Additional methods, numel and length, return the product of all dimension sizes and the size of the largest non-singleton dimension, respectively. They equal the numel and length of the entries. The size method invokes the size function, passing along parameters, on the entries. When all parameters are index objects, they convert first to numbers that address MDA dimensions. The end method, needed for subscripting operations, (,) and (,)=, with numeric subscripts, requires builtin to invoke the end function on the entries, as end is also a keyword.
The RT software uses index objects with a logical NOT operator to represent false-variant indices as subscripts, alongside true-variant ones. Whereas Table III suggests index subscripts are always required for nonzero-degree tensor expressions, the RT software enables a context-dependent index-free notation compatible with the RT algebra. For example, the inner product of an any-degree vector, x, with itself has a simpler form, x’*x, than the one shown. A tensor object always has indices, which may be specified or returned upon construction. Expressions that yield tensor results imply associated indices in predictable ways.
IV-B Index Objects
Table IV summarizes the constructor and ordinary methods, including overloaded operators, of the index class. Table V presents selected unit tests of class index. When a tensor object is constructed, its indices may be simultaneously dealt, as one or more index objects, to extra output arguments on the left-hand side (LHS) of an assignment.
TABLE IV: Index object class definition. Having few of its own methods, the index class inherits copy-by-reference, (in)equality, and concatenation operators from the handle class. Inheritance enables use of set-theoretic functions exploited by the tensor class to parse an extended Ricci notation.
Public methods (index class)
Constructor, etc.
    index (0, 1 args in), deal
Unary methods
    logical, true, false, not (~)
TABLE V: Unit tests of index objects. Each entry of an index object, which may be a vector, expresses a true-variant or false-variant state. These tests illustrate: (1) a scalar index, (2) a vector index, (3) a one-step deal, (4) a with-tensor construction, and (5) a cast-to-logical.
%% Test 1 %% Test 3
k = index; [i,j] = index(2);
assert(isscalar(k)) assert(isscalar(i))
assert(k ~= ~k) assert(isscalar(j))
assert(k == ~~k) %% Test 4
assert(true(~k) == k) sz = [100 100 10];
assert(false(k) == ~k) [A,k] = tensor(rand(sz));
%% Test 2 assert(isa(k,’index’))
k = index(2); %% Test 5
[i,j] = deal(k); k = [~index index(2)];
assert(isscalar(i)) var = logical(k);
assert(isscalar(j)) ftt = [false true true];
assert(all(k == [i j])) assert(isequal(var,ftt))
A scalar index is constructed as follows. The object has two protected properties, tilde, itself an index object, and state, a logical variable. The tilde property of the index object’s tilde property equals the index object itself. Moreover, the state of the index object and the state of its tilde property are always complementary. Upon construction, they are true and false, respectively. The constructor accomplishes this feat thanks in part to inheritance of the handle class and also via one-level recursion. A handle object resembles a void pointer in C/C++.
For a scalar index object, the logical method returns its state property (class logical) and the true or false methods return the index object itself or its tilde property (class index), whichever one has a state property equal to logical true or false, respectively. The NOT operator, overloaded by the not method, always returns the tilde property. In this way, respectively, the true-variant, false-variant, or complementary index object is obtained. As with the logical method, the true, false, and not methods work entrywise if the index object is nonscalar.
Because it inherits from class handle, class index does not need a horzcat method to enable index object concatenation into a row vector of class index. Kernel functions, like length, also work as expected. A one-argument form of the index constructor accepts a number to construct a row vector of class index with length equal to that number. The public deal method is invoked by the constructor to deal each entry of the row vector to output arguments in sequence. If there are fewer output arguments than the vector length, the last output argument is the vector remainder.
Upon construction, entries of an index object and their complements are unique, due to the handle superclass. Copy assignment, equality, and inequality operators are inherited. Entrywise equality of one index object to another, or to its complement via the entrywise NOT operator, offers a way to identify repeated indices, whether of the same variant or not, in a tensor expression. Entrywise relational operators mean that set-theoretic functions like ismember and unique, which invoke a kernel sort, may be used with the index objects. As they are currently an interpreted part of the kernel, set-theoretic functions incur abstraction penalties.
The NT software uses template metaprogramming (LibNT), character vectors (NTToolbox), and counters (LibNT and NTToolbox) to parse the multi-variant indices with underline notation of the NT algebra. In contrast, each index object of the RT software always has only two variants and, with the tensor class, the index class helps to parse overscript bar notation, which can represent superscript/subscript notation with dot spacers as required, of the RT algebra.
IV-C Unary Operations
A tensor may be subscripted with index, numeric, or other objects on the right-hand side (RHS) or LHS of an (implied) assignment via overloaded subsref and subsasgn operators, respectively. Enclosing comma-separated subscripts in parentheses, the subsref operator is considered a unary operator on the tensor the parentheses follow. The subsasgn operator is considered a binary operator where one operand is the tensor the parentheses follow, on the LHS of the assignment, and the other operand is the result of an expression, on the RHS of the assignment.
When the subsref method is invoked, either all or none of the subscripts must be index objects. In the second case, the result of the subsref expression is an equivalent subsref expression applied to the entries of the tensor. The first case invokes a simplify method or throws an error. To avoid the runtime error, there must be at least as many indices, when subscripts are concatenated, as the degree of the tensor. Extra indices address trailing singleton dimensions of the entries and increase the degree of the tensor result. This feature supports a variety of outer operations.
The protected simplify method does two things. First, for each set of indices that match, irrespective of variants, an attraction is performed. Second, for each set of attracted indices, a contraction is performed if at least one member is of the complementary variant. Attracted indices that are not contracted express unique variants that are determined and assigned within the indices property of the returned tensor. All attractions are performed via one invocation of a protected method, select, that leverages kernel support for linear indexing. All contractions are performed via one invocation of an overloaded and public sum method.
After converting index arguments to numeric dimensions, the sum method invokes a kernel sum on the entries of a tensor, passing on arguments. First, a kernel ismember tests membership of concatenated index arguments in the indices, finding dimensions of entries. With the true method, the index test is done irrespective of variants. Before invoking sum on the entries, dimensions from matched indices are increased by two, as first and second dimensions address rows and columns of entries, respectively. Unmatched index arguments address trailing singleton dimensions, which always exist, of the entries.
Though the subsasgn operator is binary, it relies on the unary permute method. The latter requires either all-index or all-numeric arguments. Before invoking a kernel permute on the entries property, passing along all arguments, the index arguments find equivalent numeric dimensions. The permute method updates the sequence of returned indices, which make up the indices property, to correspond.
In the case where all (other) arguments are index objects, when the permute method is invoked on a tensor object, a comparison is made between the indices property, the old indices, and the index arguments, the new indices. All old indices must be retained, variants unchanged. Additional new indices, which find trailing singleton dimensions, are allowed, in which case the degree increases. Each new index of a permuted tensor has the variant as specified.
Unlike subsasgn, subsref never invokes permute. A subsref operation with index subscripts ignores preexisting indices of the tensor operand. Concatenated index subscripts simply redefine indices of the returned tensor, notwithstanding contractions and attractions, which simplify entries and shorten indices, reducing degree.
The remaining unary methods of class tensor, operators included, may be divided into entrywise and pagewise groups. Entrywise unary methods, from uminus to round, apply a corresponding kernel function to the entries property. The round method, allowing a precision argument, illustrates a parameterized operation. Parameters are passed unchanged to the kernel function applied to the entries property.
Two pagewise methods, transpose and ctranspose, invoke kernel pagetranspose and pagectranspose functions on the entries of a tensor. These functions are like a for loop over third and additional dimensions of the MDA where to each page, a contiguous 2D array of the first two dimensions, a kernel transpose or ctranspose is applied. Two other kernel functions for 2D arrays, trace and diag, were likewise extended for MDAs into pagetrace and pagediag functions. They are invoked on entries of a tensor by pagewise trace and diag methods.
IV-D N-ary Operations
Kernel horzcat and vertcat functions compute column and row concatenations, respectively, for MDAs. These familiar $N$-ary operations, for vectors, matrices, and arrays of compatible dimensions, when overloaded by the tensor class as horzcat and vertcat methods, involve $N$ operands at once. Each overloaded method invokes a protected alignn method to modify each operand into a suitable MDA prior to a pagehorzcat or pagevertcat invocation that produces the entries of the initial tensor result. The alignn method also supplies the indices of the initial result. The final tensor result is produced after contractions, specified by alignn, are computed using the sum method.
The alignn method inputs $N$ and outputs $N+2$ arguments. For each tensor input only its entries, an MDA, is output. Via a kernel permute, the MDA is modified so that its dimensions align based on the set-theoretic union, in sequence, of indices irrespective of variants. A non-tensor input passes to an output only if it is a 2D array. One additional output is the set-theoretic union of indices, variants and left-to-right sequence preserved. If two variants of an index appear the leftmost is kept. The second additional output, an index vector, specifies which indices require summation, i.e., ones where two variants appeared.
To support outer concatenation, possible at non-row/column dimensions, pagehorzcat and pagevertcat invoke pagecat on the $N$ permuted MDAs from the alignn method. Corresponding non-singleton dimensions must equate in size except for the dimension at which concatenation is performed. As kernel horzcat, vertcat, and cat do not support outer concatenation, singleton dimensions in the third or higher position are expanded, using one kernel repmat per operand, to match expected sizes of corresponding dimensions after concatenation, now performable via a kernel cat.
Generalizing concatenation further, the cat method implements index concatenation. In its InferiorClasses list, the tensor class includes the index class so that the former’s cat method is invoked in a cat expression where the first argument is an index object. The cat method converts the first argument, as required, into a number that specifies the dimension of concatenation, considering the alignn step. Apart from invoking pagecat directly, passing along the dimension of concatenation, the cat method is thereafter identical to the horzcat and vertcat methods.
Like round, the isequal method features in unit test details. Implementation of isequal is similar to horzcat and vertcat. However, the isequal method instead of a tensor returns a logical, equal to the result of a kernel isequal, also $N$-ary, invoked instead of pagehorzcat or pagevertcat on the MDAs from alignn. Moreover, instead of contractions the isequal method always returns false if one operand has an index of one variant and another operand has the same of the complementary variant. This result is congruous with how the kernel isequal of a column vector and its transpose, a row vector having identical entries in the same sequence, returns false.
In theory, MATLAB could invoke a single overloaded method for an $N$-ary product expression involving a tensor. However, it invokes a sequence of binary product methods with a left-to-right precedence rule. Prior to RT software expression, this rule may require modification of an RT algebraic expression. In particular, to express an $N$-ary inner product over an index all but the rightmost operand must have the same variant, true or false, for that index, with only the rightmost operand having a complementary variant.
The NT software has two deviations from its NT algebra. To correctly compute some expressions, one or more sub-expressions have to be enclosed in parentheses and preceded by an extra operator, whose purpose is to decrement internal counters associated with underline-equivalent operators on enclosed indices. For this purpose, the tilde operator is hijacked. Thus, it cannot represent the entrywise NOT of a DenseNT, a tensor class of the NT software. These deviations represent additional weaknesses, this time from a constructivism angle, of NT underline and multi-variant index formalisms.
IV-E Binary Operations
When the subsasgn method is invoked, for a tensor operand on the LHS of an assignment, it provides parameters, namely subscripts of the LHS operand, in addition to the RHS operand. If all subscripts are index objects the assignment overwrites the LHS operand with the RHS one after its permute method is invoked so that indices are in the sequence specified by the LHS subscripts and not the RHS operand, required to be a tensor. If the RHS operand is not a tensor here, a runtime error results. If all LHS subscripts are numbers, the kernel’s subsasgn function is applied to the LHS entries with those subscripts and either the original RHS operand, if not a tensor, or its entries.
Entrywise binary methods, from plus to power in Table II, each invoke a protected method, binary, passing operands along as second and third arguments. The first argument is a handle, e.g., @plus, to an entrywise kernel function. Within binary, after it invokes a protected alignn on the operands, there are two index vectors and two MDAs. To compute initial entries of the final tensor, the entrywise kernel function is applied to the aligned MDAs. With the first index vector from alignn as initial indices, an initial tensor is produced using the two-argument constructor. Using the second index vector to specify contractions, the sum method is invoked to produce the final tensor.
Remaining binary methods, mtimes, mldivide, and mrdivide, are called pagewise because they exploit kernel pagemtimes, pagemldivide, and pagemrdivide functions through protected methods, namely mbinary, mbinary2, align2, and lattice. Each pagewise binary method invokes the mbinary method with a kernel function handle, e.g., @pagemtimes, as the first argument. For computational efficiency, the design involves a minimal number of kernel reshape and kernel permute invocations.
The pagewise binary methods exploit the lattice concept. A lattice is a degree-one matrix represented as a 3D array. The matrix may be vector or scalar, depending on dimension sizes, and there may be one page. Lattices help to realize an arbitrary inner, outer, and entrywise mtimes between two tensor operands. Each operand maps (permute-and-reshape) to a lattice, $\mathbf{A}$ or B, a lattice product executes as follows, and the lattice result, $\mathbf{C}$, maps (reshape-and-permute) to a tensor with indices as required by the RT algebra:
$\mathbf{C}_{k}$ $=\mathbf{A}_{k}\mathbf{B}_{k}\text{.}$ (98)
All outer products are represented by rows of the first lattice, $\mathbf{A}$, and columns of the second lattice, $\mathbf{B}$. All inner products are represented by columns and rows, respectively, of the first and second lattices. The common page dimension represents all entrywise products, which may execute in parallel.
Similarly, the RT software implements mldivide and mrdivide consistent with the RT algebra. Such left- or right-divisions imply a solution to a corresponding linear system, allowing inner, outer, and entrywise products. These map to and from a left- or right-division of lattices, as follows:
$\mathbf{B}_{k}$ $=\mathbf{A}^{k}\backslash\mathbf{C}_{k}\text{,}$ (99)
$\mathbf{A}_{k}$ $=\mathbf{C}_{k}/\mathbf{B}^{k}\text{.}$ (100)
As with scalar division, indices of denominator nonscalar operands have to change variants. Otherwise, the RT framework would be inconsistent in the context of linear systems. Prior to invoking mbinary, the mldivide and mrdivide methods complement indices via a not operation.
The mbinary method receives the operands of a pagewise binary method as second and third arguments. If either operand is not a tensor then it must be a 2D array to avoid a runtime error by design. With a 2D array operand the tensor result has the same indices as the tensor operand. Resulting entries equal the result of a pagewise function, specified by handle, applied to the tensor operand’s entries and the non-tensor operand. Because linear system operations are noncommutative, in general, pagewise function operands follow the original sequence received by mbinary.
When both operands of mbinary are tensor objects, the method invokes mbinary2 on them. After mbinary2 returns, mbinary gets from it the final indices of the operation, two MDAs, and two vectors that specify a kernel reshape-and-permute. To compute the final entries of the original operation, mbinary applies a pagewise function, specified via a handle, to the two MDAs, in the order received, followed by the kernel reshape-and-permute.
The mbinary2 method invokes an align2 method that depends only on indices of each operand. This results in matching and leftover indices for each operand. Additional vectors indicate whether matching indices have the same variants. After align2 is done, mbinary2 knows which dimensions of each operand’s entries correspond in inner, outer, and entrywise fashion. Using this information, a kernel permute-and-reshape is applied separately, via the lattice method, to each operand’s entries. Along with resulting MDAs, the mbinary2 method returns information to mbinary for its kernel reshape-and-permute.
Figure 7 presents representative abstraction penalties of RT software, when operands are small enough, and computational efficiencies, when operands are large enough. In examples of unary, binary, and ternary operations with compatible dimension sizes, two operands, A and B, are degree-one matrices each constructed as a tensor with the same scalar index. Other operands, C or AA and BB, are 2D or 3D arrays.
Figure 7: Abstraction penalties vs. dimension size. The interpreted RTToolbox is computationally efficient when underlying unary, binary, or ternary operations involve large enough operands. In these tests, tensor operands have the same dimension sizes, $100\times 100\times P$. Tests vary just one dimension size, $P$.
For the Figure 7 results, each operand or its entries is an array of class double with rand values. Compute times are averaged over multiple runs. Results are compared for equivalent operations done without tensor and index classes. In the binary case, the equivalent computation uses a kernel pagewise function. Although a kernel mtimes, for 2D arrays, may be applied to MDAs in pagewise fashion with one or more for loops (interpreted), the kernel pagemtimes (compiled) does the same but faster. As no kernel pagewise functions (interpreted or compiled) exist for the unary and $N$-ary cases, the comparisons use RTToolbox functions (interpreted).
The RT software expresses the RT algebra with programmatic efficiency. When abstraction penalties are small, relative to equivalent computations requiring only the kernel and novel pagewise functions, the RT software achieves computational efficiency. Because kernel ismember and unique functions, critical to parsing the RT algebra, are interpreted not compiled, they are a key source of penalties. So that the RTToolbox obeys toolbox recommendations of MATLAB Central File Exchange, it was designed without compilable MEX files, such as C/C++ implementations of kernel set-theoretic functions.
V Conclusions
This paper introduced an RT framework, the successor to a published NT framework, both of which complement a popular EMV framework. Comprising an underlying RT algebra and codesigned RT software, to model a problem and develop a solution, the proposed RT framework represents an approach for model-based imaging involving numeric tensors.
The RT algebra inherits advantages of the NT algebra, with respect to $N$-degree support, associativity, commutativity, entrywise products, linear invertibility, and other published considerations. However, due to its dual-variant index notation, as opposed to the de facto multi-variant index notation of the NT algebra, the RT algebra provides all the expressivity of the NT algebra in a simpler way. Thanks to additional outer operations, inspired by EMV broadcasting, the RT algebra is also more expressive than the NT one. Although it aligns well with the Ricci notation of geometric tensor calculus, this paper proposes the RT algebra for numeric tensor purposes.
Relevance of the RT algebra for model-based approaches to imaging is illustrated with an example, which demonstrates an asymptotically-efficient solution of an optimization problem where the SSE, a scalar function of deviant pixels in the image plane, depends on an unknown phase aberration in the pupil plane, a Fourier domain. The RT algebra helps to model and simplify the SSE and its first and second derivatives, namely a gradient and an HMF, where the complexity of computing all three proves equivalent to computing just the SSE alone. The approach makes visible diffraction rings and exoplanets near an occulted star in a simulated coronagraph image.
Thanks to the RT algebra, all nonzero-degree matrices in a gradient formulation were eliminated. Moreover, gradient and HMF expressions were reorganized to reveal (inverse) 2D DFTs, acceleratable via (inverse) 2D FFTs. Whereas the asymptotic time-and-space efficiencies were presented using MATLAB and not the RT software, the latter proved useful. A natural implementation of the RT algebra simplified a validation of intermediate steps in the underlying derivation. During the search for asymptotic efficiencies, algebraic errors were corrected via related RT software sanity checks.
The RT software is defined as the RTToolbox, comprising tensor and index classes, new pagewise functions, and unit tests, plus MATLAB. In this manner, the RT framework complements a popular EMV framework. Methods, functions, and tests presented, a subset of what could be developed, were chosen to emphasize key software requirements. Arguing that additional computational efficiency is possible, the approach taken emphasizes programmatic efficiency. Allowing the RT algebra to be naturally expressed with MATLAB, the RTToolbox is designed in a way where small abstraction penalties, related to index objects, would reduce with a MATLAB release that compiled kernel set-theoretic functions.
Compared to the NT software for MATLAB end users, comprising a C/C++ library called LibNT, M-file and MEX wrappers of the NTToolbox, plus MATLAB itself, the RT software is simpler, comprising only M-files of the RTToolbox plus MATLAB. The toolbox uses an object-oriented approach to realize a dual-variant index notation, with its index class inheriting a kernel handle class, congruous to C/C++ void pointers, for extra programmatic and computational efficiency. In contrast, while the NTToolbox has a tensor class called DenseNT, it relies on MATLAB character vectors and string processing to realize a multi-variant index notation.
Finally, the paper elaborated on unary, binary, and $N$-ary tensor operations. For example, to express an $N$-ary inner product over an index, the RT software may vary from the RT algebra. As MATLAB parses an $N$-ary product into a binary product sequence with left-to-right precedence, all but the rightmost index must express an entrywise, not inner, product. This variation is contrasted with those of the NT software, which unlike the NT algebra required some products to be enclosed in parentheses and preceded with an additional operator. Other than introducing a left-to-right precedence rule, the RT software expresses the RT algebra faithfully.
Acknowledgments
The author gratefully acknowledges editorial advice from Dan Sirbu, who suggested illustrating the RT framework with a model-based example involving DFTs, and Adam P. Harrison, whose NT framework laid a foundation for this work. He also thanks anonymous reviewers for constructive feedback that led to significant improvements, especially to clarity of the paper’s topics and goals and to its imaging example.
Code and Data
Code is shared with a permissive free software licence. For the imaging example, including its free source image, visit www.github.com/KoderKong/Fourier2D. For the RTToolbox, visit www.mathworks.com/matlabcentral/fileexchange/156174-ricci-notation-tensor-toolbox-rttoolbox.